# Intermediate mass-ratio inspirals in a dense dark-matter environment: Effects of the initial dark-matter distribution

### v1

@@ I Introduction
While there have been studies of IMRIs in DM distributions that were fit to those with angular-momentum cutoffs [50], these simulations did not include feedback onto the DM distribution. They also did not explicitly compare how different initial dark-matter distributions with a cutoff in position space versus a cutoff in angular-momentum space affect the dynamics of the IMRI. Thus, one main focus of this paper is to determine the relative amount of GW dephasing that arises in DM distributions with the more physical cutoff in angular-momentum space versus those with a cutoff in position space. To do so, we first adapt the formalism for computing feedback to the DM distribution function from dynamical-friction and secondary-accretion to be consistent with an angular-momentum cutoff. The way that we implemented this change also had an ancillary benefit of producing a large speed-up in the HaloFeedback code [51] that was used to perform the simulations. We find that for IMRIs with less extreme mass ratios, the GW dephasing between the cases with a position-space cutoff and the angular-momentum cutoff is a small fraction of the total DM-induced dephasing. This small change occurs because most of the dephasing accumulates at larger radii, where the two dark-matter distributions are very similar. For more extreme mass ratios, the GW dephasing between the inspirals in the two DM distribution functions is comparable to the total DM-induced dephasing: namely, it can be thousands of GW cycles over a four-year LISA mission.
@@ II Evolution equations for the binary and dark matter
II.1 Evolution equations for the binary
@@
Equation (4b) accounts for gravitational radiation-reaction from the quadrupole formula. In Eq. (4c), we introduced the notation $\rho_{\mathrm{DM}}(r_{2},t;v<v_{2})$ to indicate that the dynamical-friction force arises from DM particles moving more slowly than the orbital speed of the secondary ($v<v_{2}$). We also introduced the Coulomb logarithm $\log\Lambda$, which arises because it is assumed that there are minimum and maximum impact parameters of DM particles that scatter with the secondary. As in [40], we use the value $\Lambda\approx\sqrt{m_{1}/m_{2,0}}$, where $m_{2,0}$ is the initial value of $m_{2}$ (the secondary’s mass is time dependent, because of accretion).
@@
Unlike dynamical friction, secondary accretion depends on the total DM density at the location of the secondary, not just of the particles moving more slowly than the secondary.
@@ II.2 Dark-matter distribution function and density
II.2.1 Dark-matter distribution with a position-space cutoff
@@ II.3 Dynamical-friction feedback with an angular-momentum cutoff
$\times\delta\boldsymbol{(}\Delta\mathcal{E}(b)-\Delta\mathcal{E}\boldsymbol{)}\,.$ (28)
The change in energy $\Delta\mathcal{E}$ is parameterized in terms of the impact parameter $b$ of a DM particle scattering with the secondary. The total scattering rate at a given energy is obtained by integrating over all allowed changes in energy $\Delta\mathcal{E}$,
@@
We also introduced the notation
@@ III Special cases of the minimum angular momentum
III.1 No angular-momentum cutoff
@@ III.2 Properties of the dark matter distribution for a nonzero angular-momentum cutoff
Figure 1: Dark-matter density with different angular-momentum cutoffs. In both panels, we show the density calculated from the distribution function given by Eq. (12). We consider here a primary mass $m_{1}=10^{5}\,\mathrm{M}_{\odot}$, $\rho_{\mathrm{sp}}=200\,\mathrm{M}_{\odot}/\mathrm{pc}^{3}$ and $\gamma_{\mathrm{sp}}=7/3$. Left: The analytic power law density ($j_{\mathrm{min}}=0$) in dotted black, the fully Newtonian result ($j_{\mathrm{min}}=4$) in dashed light-gray, the fully relativistic result in dash-dotted orange, and the modified Newtonian result ($j_{\mathrm{min}}=\sqrt{8}$) in solid blue. Right: The latter two densities, on linear scale on the vertical axis. We plot a dotted, black, vertical line at $4\,R_{\mathrm{s}}=r_{\mathcal{J}}$ for reference. The figure is discussed in more detail in the text of Sec. III.2.
@@ III.3 Feedback formalism with a minimum angular momentum
We also briefly comment on dynamical heating, in which particles moving more quickly than the secondary can transfer energy to the secondary. This process would partially counter dynamical friction, and it could, in principle, increase a given particle’s relative specific energy in excess of the maximum $c^{2}/(2j_{\mathrm{min}}^{2})$, leading it to be captured by the primary. We do not consider dynamical heating in this paper, given that it is suppressed relative to dynamical friction by a factor of the dark matter particle mass over the secondary’s mass, which would be small for any particle dark matter. Though note that Ref. [42] has taken into account the effects from faster moving particles in Chandrasekhar’s formula for IMRIs.
@@ IV.1 Simulations and initial conditions
We simulate the inspiral of five binaries with different mass ratios using a version of the HaloFeedback code [51] which was modified to have a nonzero $j_{\mathrm{min}}$. This code is named HaloFeedbackAcc and is available on GitLab [56]. We initialize the mass of the secondary to be $m_{2}=10\,\mathrm{M}_{\odot}$, and we vary the primary mass. The five values we choose are $m_{1}=$ $10^{3}\,\mathrm{M}_{\odot}$, $m_{1}=$ $3\times 10^{3}\,\mathrm{M}_{\odot}$, $m_{1}=10^{4}\,\mathrm{M}_{\odot}$, $m_{1}=3\times 10^{4}\,\mathrm{M}_{\odot}$ and $m_{1}=10^{5}\,\mathrm{M}_{\odot}$. For each simulation, we initialize the dark-matter distribution function to that given in Eq. (12), with $\rho_{\mathrm{sp}}$ = $200~\mathrm{M}_{\odot}$ and $\gamma_{\mathrm{sp}}$ = $7/3$. We will call an inspiral in such an initial dark-matter distribution a “first-generation” inspiral. The reason for this naming is that the density profile is consistent with a primary black hole that grew adiabatically in mass and did not undergo a previous IMRI inspiral (which would have modified the distribution of dark matter through feedback).
@@ IV.2 Gravitational-wave dephasing
Table 1: Number of gravitational-wave cycles, dephasing, and characteristic masses for inspirals with different $\boldsymbol{j_{\mathrm{min}}}$. We denote $N_{\mathrm{cycles}}^{\mathrm{A}}$ as the number of GW cycles between a given starting radius and ISCO for a simulation corresponding to case $\mathrm{A}$, where $\mathrm{A=1}$ is the case with no cutoff in angular momentum for a first-generation inspiral, and $\mathrm{A=2_{I}}$ is the case with the angular momentum cutoff with $j_{\mathrm{min}}=\sqrt{8}$ for a first-generation inspiral. The dephasing $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-A)}}$ is the difference in cycles between a vacuum inspiral and case $\mathrm{A}$ (both of which are computed from the starting radius of $r_{\mathrm{2,4y}}^{\mathrm{A}}$ described in the text of Sec. IV.2). For the vacuum system, we use $m_{2}=10\,\mathrm{M}_{\odot}+\delta m_{\mathrm{2,4y}}^{\mathrm{A}}$, and list $\delta m^{\mathrm{A}}_{\mathrm{2,4y}}$ (again for the cases $\mathrm{A=1}$ and $\mathrm{A=2_{I}}$). The dephasing is computed for the last four years of inspiral before the secondary reaches the ISCO, using the initial separations $r_{\mathrm{2,4y}}^{\mathrm{A}}$. As summarized in Sec. IV.1, the initial DM distribution is characterized by $\gamma_{\mathrm{sp}}=7/3$, and $\rho_{\mathrm{sp}}=200\,\mathrm{M}_{\odot}/\mathrm{pc}^{3}$.
$m_{1}[\mathrm{M}_{\odot}]$ $N_{\mathrm{cycles}}^{\mathrm{1}}$ $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-1)}}$ $\delta m_{\mathrm{2,4y}}^{\mathrm{1}}[\mathrm{M}_{\odot}]$ $N_{\mathrm{cycles}}^{\mathrm{2_{I}}}$ $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-2_{I})}}$ $\delta m_{\mathrm{2,4y}}^{\mathrm{2_{I}}}[\mathrm{M}_{\odot}]$
@@
Dynamical friction and secondary accretion cause the binary to inspiral more rapidly than without these effects in vacuum. This leads to a dephasing of the gravitational-wave signal from an analogous vacuum inspiral, given that the binary undergoes fewer orbital cycles before reaching the ISCO. One can compute the number of cycles from
$N_{\mathrm{cycles}}=\frac{1}{\pi}\int_{r_{2,\mathrm{i}}}^{r_{\mathrm{ISCO}}}\!\Omega\,\dot{r}_{2}^{-1}dr_{2}\,.$ (51)
@@
$\Delta N^{\mathrm{(A-B)}}_{\mathrm{cycles}}=N^{\mathrm{A}}_{\mathrm{cycles}}-N^{\mathrm{B}}_{\mathrm{cycles}}$ (52)
for two cases $\mathrm{A}$ and $\mathrm{B}$. As in [48], we choose $\mathrm{A}=0$ to denote the vacuum case, $1$ to the case with dark matter and with $j_{\mathrm{min}}=0$, and $\mathrm{2_{I}}$ to the be with dark matter, with $j_{\mathrm{min}}=\sqrt{8}$ and for a first-generation inspiral.
Because the secondary accretes mass throughout the inspiral, we must make a choice of an appropriate mass $m_{2}$ and initial orbital separation $r_{2,\mathrm{i}}$ for an equivalent vacuum system against which we compute the dephasing. In Ref. [48], because the mass $m_{2}$ was fixed in the evolution equations, it was most natural to compare against a vacuum system with $m_{2}=10\,\mathrm{M}_{\odot}$ and $r_{2,\mathrm{i}}=r_{\mathrm{2,4y}}$ (the appropriate radius for a system with a secondary mass of $m_{2}=10\,\mathrm{M}_{\odot}$). Here, we will compare with a vacuum system with $m_{2}=m_{\mathrm{2,4y}}^{\mathrm{A}}$, where $m_{\mathrm{2,4y}}^{\mathrm{A}}$ is the mass of the secondary for the case $\mathrm{A}$ at the radius $r_{\mathrm{2,4y}}^{\mathrm{A}}$ for a vacuum system with the same secondary mass as $m_{\mathrm{2,4y}}^{\mathrm{A}}$. Because this is an implicit relation, since we do not know $m_{\mathrm{2,4y}}^{\mathrm{A}}$ a priori, we determine $m_{\mathrm{2,4y}}^{\mathrm{A}}$ and $r_{2,\mathrm{i}}^{\mathrm{A}}$ using a simple two-step procedure. First, we determine the value of $m_{2}=m_{\mathrm{2,4y}}^{\mathrm{A}}$ at the radius $r_{\mathrm{2,4y}}$ corresponding to a vacuum system with $m_{2}=10\,\mathrm{M}_{\odot}$. Second, we determine $r_{\mathrm{2,4y}}^{\mathrm{A}}$ for the vacuum system using the value $m_{2}=m_{\mathrm{2,4y}}^{\mathrm{A}}$ from the first step, and choose $r_{2,\mathrm{i}}^{\mathrm{A}}=r_{\mathrm{2,4y}}^{\mathrm{A}}$.
We present in Table 1 the cycles and dephasing against vacuum for the five simulations described in Sec. IV.1. We also show a change in mass, for which we will introduce the notation
$m_{\mathrm{2,4y}}^{\mathrm{A}}=m_{2,0}+\delta m_{\mathrm{2,4y}}^{\mathrm{A}}\,$, (53)

### v2

@@ I Introduction
I.1 Dark-matter effects on the gravitational-wave emission from black-hole binaries
@@
I.2 Discussion of the effects of the initial dark-matter distribution
@@
While there have been studies of IMRIs in DM distributions that were fit to those with angular-momentum cutoffs [59, 50], these simulations did not include feedback onto the DM distribution. They also did not explicitly compare how different initial dark-matter distributions with a cutoff in position space versus a cutoff in angular-momentum space affect the dynamics of the IMRI. Thus, one main focus of this paper is to determine the relative amount of GW dephasing that arises in DM distributions with the more physical cutoff in angular-momentum space versus those with a cutoff in position space. To do so, we first adapt the formalism for computing feedback to the DM distribution function from dynamical-friction and secondary-accretion to be consistent with an angular-momentum cutoff. The way that we implemented this change also had an ancillary benefit of producing a large speed-up in the HaloFeedback code [44] that was used to perform the simulations. We find that for IMRIs with less extreme mass ratios, the GW dephasing between the cases with a position-space cutoff and the angular-momentum cutoff is a small fraction of the total DM-induced dephasing. This small change occurs because most of the dephasing accumulates at larger radii, where the two dark-matter distributions are very similar. For more extreme mass ratios, the GW dephasing between the inspirals in the two DM distribution functions is comparable to the total DM-induced dephasing: namely, it can be thousands of GW cycles over a four-year LISA mission.
@@ II Evolution equations for the binary and dark matter
II.1 Review of the evolution equations for the binary
@@
Equation (4b) accounts for gravitational radiation-reaction from the quadrupole formula. In Eq. (4c), we introduced the notation $\rho_{\mathrm{DM}}(r_{2},t;v<v_{2})$ to indicate that the dynamical-friction force arises from DM particles moving more slowly than the orbital speed of the secondary ($v<v_{2}$). [^4] We restrict to $v<v_{2}$, because as is described in detail in [21], the more rapidly moving particles contribute to dynamical heating. However, dynamical heating is smaller than dynamical friction by a factor of the ratio of the DM particle mass to the mass of the secondary black hole (i.e., $m_{\mathrm{DM}}/m_{2}$). For particle DM, as we consider in this paper, this ratio is minuscule. Dynamical heating could be relevant if the distribution function were such that the ratio of the density of particles moving faster than the orbital speed to that of the more slowly moving particles were of the order of $m_{2}/m_{\mathrm{DM}}$. However, for the DM spike density that we consider in this paper, the ratio of these densities is of order one (see, e.g., [46, 53]); thus, dynamical heating will be negligible, and we do not consider it. While there have been results in the literature that state that dynamical heating can be significant (e.g., [30]), Ref. [30], for example, neglects the suppression of dynamical heating by $m_{\mathrm{DM}}/m_{2}$, thereby implicitly assuming the ratio is order one. This could be appropriate for dark matter formed from primordial black holes, but this is not the scenario which we consider in this paper. We also introduced the Coulomb logarithm $\log\Lambda$, which arises because it is assumed that there are minimum and maximum impact parameters of DM particles that scatter with the secondary. As in [46], we use the value $\Lambda\approx\sqrt{m_{1}/m_{2,0}}$, where $m_{2,0}$ is the initial value of $m_{2}$ (the secondary’s mass is time dependent, because of accretion).
@@
Equation (5) assumes the typical relative velocity between the secondary and a DM particle is the secondary’s velocity $v_{2}$ (see, e.g., [53] and references therein). Under this assumption, the cross section can be obtained from a calculation of the critical impact parameter for a Schwarzschild black hole to capture particles on geodesics with an asymptotic velocity of $v_{2}$ (see, e.g., Exercise 25.22 of [49]). Thus, Eq. (5) is appropriate for quantifying the accretion of locally isotropic distributions of collisionless particle dark matter; it is not obviously applicable to more complex accretion scenarios.
@@
Secondary accretion depends on the total DM density at the location of the secondary, unlike dynamical friction which only depends on the density of DM particles moving more slowly than the secondary [see the $v<v_{2}$ in the density in Eq. (4c)].
@@ II.2 Dark-matter distribution function and density
II.2.1 Review of the dark-matter distribution with a position-space cutoff
@@ II.3 Dynamical-friction feedback with an angular-momentum cutoff
$\times\delta\boldsymbol{(}\Delta\mathcal{E}(b)-\Delta\mathcal{E}\boldsymbol{)}\,$, (28)
where $T_{2}=2\pi r_{2}/v_{2}$ is the orbital period of the secondary. The change in energy $\Delta\mathcal{E}$ is parameterized in terms of the impact parameter $b$ of a DM particle scattering with the secondary. The total scattering rate at a given energy is obtained by integrating over all allowed changes in energy $\Delta\mathcal{E}$,
@@
where $T_{2}$ is the orbital period of the secondary. We also introduced the notation
@@ III Special cases of the minimum angular momentum
III.1 Review of having no angular-momentum cutoff
@@ III.2 Properties of the dark matter distribution for a nonzero angular-momentum cutoff
Figure 1: Dark-matter density with different angular-momentum cutoffs. In both panels, we show the density calculated from the distribution function given by Eq. (12). We consider here a primary mass $m_{1}=10^{5}\,\mathrm{M}_{\odot}$, $\rho_{\mathrm{sp}}=200\,\mathrm{M}_{\odot}/\mathrm{pc}^{3}$ and $\gamma_{\mathrm{sp}}=7/3$. We plot a dotted, black, vertical line at $r_{\mathcal{J}}=4\,R_{\mathrm{s}}=8Gm_{1}/c^{2}$ for reference. Left: The analytic power law density ($j_{\mathrm{min}}=0$) in dotted black, the fully Newtonian result ($j_{\mathrm{min}}=4$) in dashed light-gray, the fully relativistic result in dash-dotted orange, and the modified Newtonian result ($j_{\mathrm{min}}=\sqrt{8}$) in solid blue. Right: The latter two densities, on linear scale on the vertical axis. The figure is discussed in more detail in the text of Sec. III.2.
@@ III.3 Feedback formalism with a minimum angular momentum
We briefly comment on dynamical heating, as discussed in Footnote 4, in the context of a nonzero angular-momentum cutoff. This process could increase a DM particle’s relative specific energy in excess of the maximum $c^{2}/(2j_{\mathrm{min}}^{2})$, leading it to be captured by the primary. However, we argued in Footnote 4 that dynamical heating is negligible for particle dark matter and for the distribution functions considered in this paper. Thus, we do not consider such an effect, which is consistent with our (well-justified) neglect of dynamical heating in the dynamics of the binary.
@@ IV.1 Simulations and initial conditions
We simulate the inspiral of five binaries with different mass ratios using a version of the HaloFeedback code [44] which was modified to have a nonzero $j_{\mathrm{min}}$. This code is named HaloFeedbackAcc and is available on GitLab [63]. We initialize the mass of the secondary to be $m_{2}=10\,\mathrm{M}_{\odot}$, and we vary the primary mass. The five values we choose are $m_{1}=$ $10^{3}\,\mathrm{M}_{\odot}$, $m_{1}=$ $3\times 10^{3}\,\mathrm{M}_{\odot}$, $m_{1}=10^{4}\,\mathrm{M}_{\odot}$, $m_{1}=3\times 10^{4}\,\mathrm{M}_{\odot}$ and $m_{1}=10^{5}\,\mathrm{M}_{\odot}$. For each simulation, we initialize the dark-matter distribution function to that given in Eq. (12), with $\rho_{\mathrm{sp}}$ = $200~\mathrm{M}_{\odot}/$pc3 and $\gamma_{\mathrm{sp}}$ = 7/37/3.999These values of $\rho_{\mathrm{sp}}$ and $\gamma_{\mathrm{sp}}$ are comparable to those used in the study [32], which have been used as a benchmark system in several subsequent works on this subject (see, e.g., [67, 33, 46, 28, 53, 65]). While it would be of interest to study a wider range of this parameter space and to investigate any correlations in this parameter space on the gravitational-wave observables, we leave such studies to future work. We will call an inspiral in such an initial dark-matter distribution a “first-generation” inspiral. The reason for this naming is that the density profile is consistent with a primary black hole that grew adiabatically in mass and did not undergo a previous IMRI inspiral (which would have modified the distribution of dark matter through feedback).
@@
We have made the data files for the DM density profiles from each simulation available to download from [64]. These data files contain the densities when the secondary is at $3r_{\mathrm{2,4y}}$, at $r_{\mathrm{2,4y}}$, and at ISCO. Data for $j_{\mathrm{min}}=0$ (first-generation inspirals) and $j_{\mathrm{min}}=\sqrt{8}$ (first- and second-generation inspirals) are available at [64].
@@ IV.2 Gravitational-wave dephasing
IV.2.1 Definitions and subtleties in computing the dephasing
Table 1: Number of gravitational-wave cycles, dephasing, and characteristic masses for inspirals with different $\boldsymbol{j_{\mathrm{min}}}$. We denote $N_{\mathrm{cycles}}^{\mathrm{(A)}}$ [see Eq. (51)] as the number of GW cycles between a given starting radius and ISCO for a simulation corresponding to case $\mathrm{A}$, where $\mathrm{A=1}$ is the case with no cutoff in angular momentum for a first-generation inspiral, and $\mathrm{A=2_{I}}$ is the case with the angular momentum cutoff with $j_{\mathrm{min}}=\sqrt{8}$ for a first-generation inspiral. The dephasing $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-A)}}$ [see Eq. (52)] is the difference in cycles between a vacuum inspiral and case $\mathrm{A}$ (both of which are computed from the starting radius of $r_{\mathrm{2,4y}}^{\mathrm{(A)}}$ described in the text of Sec. IV.2). For the vacuum system, we use $m_{2}=10\,\mathrm{M}_{\odot}+\delta m_{\mathrm{2,4y}}^{\mathrm{(A)}}$ [see Eq. (53)], and list $\delta m^{\mathrm{(A)}}_{\mathrm{2,4y}}$ (again for the cases $\mathrm{A=1}$ and $\mathrm{A=2_{I}}$). The dephasing is computed for the last four years of inspiral before the secondary reaches the ISCO, using the initial separations $r_{\mathrm{2,4y}}^{\mathrm{(A)}}$. As summarized in Sec. IV.1, the initial DM distribution is characterized by $\gamma_{\mathrm{sp}}=7/3$, and $\rho_{\mathrm{sp}}=200\,\mathrm{M}_{\odot}/\mathrm{pc}^{3}$.
$m_{1}[\mathrm{M}_{\odot}]$ $N_{\mathrm{cycles}}^{(1)}$ $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-1)}}$ $\delta m_{\mathrm{2,4y}}^{(1)}[\mathrm{M}_{\odot}]$ $N_{\mathrm{cycles}}^{\mathrm{(2_{I})}}$ $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-2_{I})}}$ $\delta m_{\mathrm{2,4y}}^{\mathrm{(2_{I})}}[\mathrm{M}_{\odot}]$
@@ IV.2.1 Definitions and subtleties in computing the dephasing
Dynamical friction and secondary accretion cause the binary to inspiral more rapidly than without these effects in vacuum. This leads to a dephasing of the gravitational-wave signal from an analogous vacuum inspiral, given that the binary undergoes fewer orbital cycles before reaching the ISCO. One can compute the number of cycles in the time domain for a case $\mathrm{A}$ by
$N_{\mathrm{cycles}}^{\mathrm{(A)}}=\frac{1}{\pi}\int_{t_{\mathrm{i}}}^{t^{\mathrm{(A)}}_{\mathrm{ISCO}}}\!\Omega\,dt\,.$ (51)
We have used the convention that the time is the same value $t=t_{\mathrm{i}}$ for all cases when the secondary is at the initial radius, and the time at which the particle reaches the ISCO in a case $\mathrm{A}$ is $t^{\mathrm{(A)}}_{\mathrm{ISCO}}$. As in [53], we choose $\mathrm{A}=0$ to denote the vacuum case; here we choose $1$ to the case with dark matter and with $j_{\mathrm{min}}=0$ and $\mathrm{2_{I}}$ to the be with dark matter, with $j_{\mathrm{min}}=\sqrt{8}$ and for a first-generation inspiral. [^10] In [53], the case $\mathrm{A}=1$ corresponded to a hypothetical, unphysical scenario in which dynamical friction and feedback occurs, but accretion does not occur. We do not consider such a case in this paper, because Ref. [53] demonstrated that both dynamical friction and accretion (with both types of feedback) should be treated to accurately represent the DM effects on the orbital dynamics, the DM distribution, and the gravitational waves emitted from these systems. Thus, we will compare simulations with all the relevant DM effects against vacuum systems without any DM effects in this paper.
@@
$\Delta N^{\mathrm{(A-B)}}_{\mathrm{cycles}}=N^{\mathrm{(A)}}_{\mathrm{cycles}}-N^{\mathrm{(B)}}_{\mathrm{cycles}}$ (52)
for two cases $\mathrm{A}$ and $\mathrm{B}$. The dephasing gives a simple qualitative assessment of the extent to which two waveforms agree or disagree. More quantitative assessments of distinguishability require computing Bayes factors or evidence ratios between different models.
Because the secondary accretes mass throughout the inspiral, there are some subtleties that arise when computing the dephasing of a system with dark matter against an “equivalent” vacuum system, which we now discuss.
First, note that in Ref. [53] such subtleties did not arise, because $m_{2}$ was held fixed in the evolution equations for the secondary, so that the chirp mass remained constant during the inspiral. However, a fractional change in $m_{2}$ of order $10^{-3}$ produces a comparable fractional change in the chirp mass (and thus also a comparable fractional change in the number of gravitational-wave cycles). For systems that undergo of order $10^{6}$ cycles, this is of order $10^{3}$ cycles. Thus, it is important to account for the changing secondary mass during the inspiral in the equations of motion, so we do not repeat the approach of [53] here.
Next, as discussed in Sec. IV.1, to obtain appropriate initial conditions at the radius $r_{\mathrm{2,4y}}$, we begin evolving the binary at a separation of $3r_{\mathrm{2,4y}}$. As it evolves between these radii, the mass changes from the initial mass $m_{2,0}=10\,\mathrm{M}_{\odot}$ by an amount which we denote $\delta m_{\mathrm{2,4y}}$. Because we aim to quantify the effects of dark matter on the inspiral from a most closely equivalent vacuum system during the final four years of the inspiral, we must consider what is a reasonable choice of “most closely equivalent.” A natural definition is a vacuum IMRI that has the same mass $m_{2}+\delta m_{\mathrm{2,4y}}$ at the radius $r_{\mathrm{2,4y}}$ as the IMRI with dark matter. Such a vacuum system will have its gravitational radiation reaction be equivalent at the radius $r_{\mathrm{2,4y}}$ to that of the IMRI with dark matter, so any changes to the inspiral from that radius inward will arise from DM effects only.
Finally, there are yet more subtle aspects to the comparison against an “equivalent vacuum IMRI” that we now discuss: (i) The change in the mass $\delta m_{\mathrm{2,4y}}$ should, in fact, be labeled $\delta m_{\mathrm{2,4y}}^{\mathrm{(A)}}$, because it will depend on which case $\mathrm{A}$ we are considering. (ii) We do not know the value of $m_{\mathrm{2,4y}}^{\mathrm{(A)}}$ a priori; it is determined after running a simulation from the radius $3r_{\mathrm{2,4y}}$ to $r_{\mathrm{2,4y}}$ (iii) We determine $r_{\mathrm{2,4y}}$ from the equivalent vacuum binary with a secondary mass that we denote by $m^{\mathrm{(A)}}_{\mathrm{2,4y}}$, which is defined from
$m_{\mathrm{2,4y}}^{\mathrm{(A)}}=m_{2,0}+\delta m_{\mathrm{2,4y}}^{\mathrm{(A)}}$ (53)
(with $m_{2,0}=10\,\mathrm{M}_{\odot}$, as before). This implies that there is also an a priori unknown radius $r_{\mathrm{2,4y}}^{\mathrm{(A)}}$, the initial radius to inspiral to the ISCO in four years of the equivalent vacuum IMRI with mass $m_{\mathrm{2,4y}}^{\mathrm{(A)}}$.
We determine $m_{\mathrm{2,4y}}^{\mathrm{(A)}}$ and $r_{\mathrm{2,4y}}^{\mathrm{(A)}}$ approximately by first computing $m_{\mathrm{2,4y}}^{\mathrm{(A)}}$ at the radius $r_{\mathrm{2,4y}}$ for a vacuum system with $m_{2}=10\,\mathrm{M}_{\odot}$, and then we determine $r_{\mathrm{2,4y}}^{\mathrm{(A)}}$ for the vacuum system using the value $m_{\mathrm{2,4y}}^{\mathrm{(A)}}$ from the first step. Thus, our equivalent vacuum IMRI that we use for computing dephasing numbers and curves will be vacuum IMRI with mass $m_{\mathrm{2,4y}}^{\mathrm{(A)}}$ starting from a radius $r_{\mathrm{2,4y}}^{\mathrm{(A)}}$, with $m_{\mathrm{2,4y}}^{\mathrm{(A)}}$ and $r_{\mathrm{2,4y}}^{\mathrm{(A)}}$ computed via the approximate procedure defined in this paragraph.
IV.2.2 Discussion of dephasing results